OpenStudy (anonymous):

Find all the square roots of -196.

OpenStudy (solomonzelman):

\[-196 = 4 \times 7 \times 7 \times( -1)\] (group factored) NOW,\[=2^2 \times 7^2 \times i^2=(14i)^2\]

OpenStudy (solomonzelman):

So your roots would be \[14i ~~~~~~~~and~~~~~~~~~~~~~~-14i\]